OpenStudy (anonymous):

Find f(x) and g(x) so that the function can be described as y = f(g(x)). y=2/(x^2) +9

OpenStudy (dumbcow):

g(x) = x^2 f(x) = 2/(x+9)

OpenStudy (anonymous):

@keeponbleeding what are the choices?

OpenStudy (anonymous):

wait how did that happen?

OpenStudy (anonymous):

no multiple choices

OpenStudy (anonymous):

is that y? \[y = \frac{ 2 }{ x^2 }+9\]

OpenStudy (anonymous):

yes

OpenStudy (dumbcow):

y = f(g(x)) means that g(x) is input for variable of f(x) say f(x) = 2/? + 9 , this is general form so ? must be x^2 thus g(x) = x^2

OpenStudy (anonymous):

thanks! got it